# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

### v1

@@ C.3 Bounds for large step-sizes $\gamma L>1$
Theorem C.9 (Bounds for SGD. Convex case, large step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu=0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(1,2)$. Then, for every $T\geq 1$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}\delta}{2\gamma(2-\gamma L)T}+\frac{\gamma\sigma_{*}^{2}\bar{e}_{T}^{\prime}}{2(2-\gamma L)^{3}},$
where $\delta=1-(1-\gamma L)^{2T}\in(0,1)$, and $\bar{e}_{T}^{\prime}$ is defined in (6) and grows exponentially with $T$. If moreover $f$ is convex, the above bound holds true for $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$, where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
We simply combine Proposition C.8 with Theorem B.4 and Lemma B.2. ∎
We may moreover show that the bias term in Theorem C.9 is tight, in the sense that no Lyapunov proof using energies of our form may improve upon the bias term.
Proposition C.10.
Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\dots,T-1$. If $\gamma L>1$ and $a_{0}=1$, then necessarily
$\rho\leq\frac{2\gamma(2-\gamma L)}{1-(1-\gamma L)^{2T}}.$
Proof.
We consider $d=1$ and $m=2$, and let $f\colon\mathbb{R}\to\mathbb{R}$ be defined through
One can easily compute that
We know that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ for all $t=0,\ldots,T-1$ and for all $x_{0}\in\mathbb{R}$. We select $x_{0}$ sufficiently large such that $x_{t}$ is sufficiently large (to be defined) for all $t=0,\ldots,T-1$. Expanding the Lyapunov decrease and rearranging yields
As $x_{t}$ is sufficiently large, and the right-hand side is constant, $a_{t+1}(1-\gamma L)^{2}-a_{t}+\frac{\rho L}{2}$ must be nonnegative, or equivalently
$a_{t+1}\leq\frac{1}{(1-\gamma L)^{2}}a_{t}-\frac{L}{2(1-\gamma L)^{2}}\rho.$
To simplify the notations, we note $\tau=(1-\gamma L)^{2}$, such that $a_{t+1}\leq\tau^{-1}a_{t}-\tau^{-1}\tfrac{\rho L}{2}$. By induction, we deduce that $a_{T}\leq\tau^{-T}a_{0}-\tfrac{\rho L}{2}\sum_{t=1}^{T}\tau^{-t}$. Because $a_{T}\geq 0$, we obtain that
$1=a_{0}\geq\frac{\rho L}{2}\sum\limits_{t=1}^{T}\tau^{T-t}=\frac{\rho L}{2}\sum\limits_{t=0}^{T-1}\tau^{T-t-1}\iff\rho\leq\frac{2a_{0}}{L\sum_{t=0}^{T-1}(1-\gamma L)^{2t}}=\frac{2\gamma(2-\gamma L)}{1-(1-\gamma L)^{2T}},$
concluding the proof. ∎
C.4 Detailed comparison to previous works
For comparative purposes, the results cited from previous works below have been adapted to align with our setting, such as by considering constant step-sizes where non-constant ones were originally used, or by adjusting the more general settings to match ours.
To the best of our knowledge, the analysis of SGD for convex functions with constant step-sizes $\gamma L\geq 1$ has not been studied previously. In the subsections below, we compare our Theorem 2.3, which addresses the case $\gamma L<1$, to the existing literature.
Comparison to gower_sgd_2021.
In (gower_sgd_2021, Theorem D.6), the authors analyzed the standard SGD algorithm under assumptions comparable to ours, while also accommodating for variable step-sizes. When adapted to our constant step-size setting, their result for $\gamma\in(0,\frac{1}{2L})$ boils down to
$\min_{k=0,\ldots,T-1}\mathbb{E}[f(x_{t})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma(1-2\gamma L)T}+\frac{\gamma}{1-2\gamma L}\sigma_{*}^{2}.$
One sees that this concurrent bound diverges when $\gamma$ gets close to $\tfrac{1}{2L}$, on both the bias and variance terms. This comparison must of course take into account the fact that the authors are able to provide bounds under weaker assumptions than our convexity and smoothness assumptions. They indeed assume $f$ to be smooth (instead of each $f_{i}$) and relax the functions $f_{i}$ to be quasar-convex, which is a consequence of our expected cocoercivity assumption, see Remark B.5.
Comparison to garrigos_handbook_2024.
Recently, in (garrigos_handbook_2024, Theorem 5.5), the finite-sum minimization problem for $L$-smooth convex functions under assumptions identical to ours was considered. For $\gamma\in(0,\frac{1}{4L}]$, they derived the bound
$\mathbb{E}[f(\overline{x}_{T})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{\gamma T}+2\gamma\cdot\sigma_{*}^{2}.$
This result is encompassed as a consequence of Theorem 2.3.
Furthermore, their analysis can easily be extended. For any $\varepsilon>0$ and $\gamma\in(0,\frac{1}{(1+\varepsilon)L})$, it holds that
$\mathbb{E}[f(\overline{x}_{T})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T(1-(1+\varepsilon)\gamma L)}+\frac{(1+\varepsilon^{-1})\gamma}{2(1-(1+\varepsilon)\gamma L)}\cdot\sigma_{*}^{2}.$
This generalized bound is also recovered by Theorem 2.3 for any $\varepsilon>0$. In the interpolation setting, allowing $\varepsilon\to 0$, Theorem 2.3 remains strictly stronger.
Comparison to taylor_stochastic_2019.
In (taylor_stochastic_2019, Theorem 15), the authors investigate a primal averaging variant of SGD, originally with variable step-sizes. With constant step-sizes, their algorithm generates an auxiliary sequence $(y_{k})$ and follows the update rule
$y_{k+1}=\frac{k}{k+1}y_{k}+\frac{1}{k+1}x_{k},\quad x_{k+1}=x_{k}-\gamma\nabla f_{i_{k}}(y_{k+1}),$
where $i_{t}\in\mathcal{I}$ is sampled i.i.d. from the distribution $\mathcal{D}$.
For $\gamma\in(0,\frac{1}{L})$, they establish the bound
$\mathbb{E}[f(y_{T})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma}{2(1-\gamma L)}\cdot\sigma_{*}^{2}.$
We observe that Theorem 2.3 provides a tighter bound, achieved without the use of the primal averaging scheme. It is worth noting that asymptotically, as $N\to\infty$, the two bounds are equivalent.
The authors also studied the interpolation setting, when $\sigma_{*}^{2}=0$, for their primal averaging scheme over a wider range of step-sizes (taylor_stochastic_2019, Theorem 8). For $\gamma\in(0,\tfrac{2}{L})$, they obtained
$\mathbb{E}\mathopen{}\left[{f(y_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T(1-[\gamma L-1]_{+})},$
which is also weaker than Theorem 2.3, Corollary 2.6 and Theorem 2.7.
Regarding the primal averaging scheme, the comparison above suggests that for constant step-sizes in this setting, it does not yield faster convergence compared to standard SGD. This observation contrasts with the recent comparison of primal averaging schemes with acceleration schemes (defazio_momentum_2021).
We note that the authors also obtain rates on the vanilla SGD algorithm, where they however assume a uniformly bounded gradient variance $\sup_{x}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)-\nabla f(x)\|^{2}}\right]\mathclose{}\leq\sigma^{2}$. Specifically, (taylor_stochastic_2019, Theorem 5) obtains a last iterate convergence rate for $\gamma L\in(0,1]$, given by
$\mathbb{E}\mathopen{}\left[{f(x_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\gamma v_{T}\sigma^{2}\text{ with }v_{T}\simeq T.$
They additionally obtain in (taylor_stochastic_2019, Theorem 6) a bound on the Cesaro average for $\gamma L\in(0,\tfrac{1+\sqrt{5}}{2}]$, given by
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\gamma\frac{1+L}{2}\sigma^{2}.$
Appendix D Proofs in the smooth strongly convex setting
@@
As in the previous section, we derive parameters that satisfy the sufficient conditions of Theorem B.4 in order to derive upper bounds for SGD, this time in the strongly convex setting. We again note that all propositions could simply be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
D.1 Bounds for any step-size
@@
Because it seems impossible to obtain a bound with tight bias term for optimal step-sizes (see Sections F.3 and F.4), we provide here a more relaxed bound, where the bias term is $\varepsilon$ away from being tight, and subsequently the variance term remains bounded for every step-size $\gamma L\in(0,2)$.
@@
Proposition D.1 (Lyapunov parameters. Strongly convex case, any step-sizes, $L>\mu$).
@@
Let $\gamma L\in(0,2)$, with $L>\mu$. Let $\varepsilon\geq 0$ and $\gamma_{opt}=\tfrac{2}{\mu+L}$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{opt}$. Consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
@@
  • $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{opt}^{2}+\varepsilon$, with
    $\phi_{opt}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{opt}),$
@@
    $\omega=\frac{\gamma\phi_{opt}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}\quad\text{ with }\quad\delta=|\gamma-\gamma_{opt}|,$
@@
Then the sufficient Lyapunov conditions of Theorem B.4 are verified. Moreover, we have
@@
Proof.
We start the proof by exploiting some insights from our numerical analysis. First, we observe that we can impose $\rho=0$ and $\alpha_{t}=\beta_{t}$ without changing anything to the empirical best bias term. We moreover infer that ${a_{t}}/{a_{t+1}}\in(0,1)$ is constant, so let us denote this ratio by $\phi^{2}$. Since $\phi_{\text{opt}}^{2}$ is the optimal rate for the deterministic gradient descent in the strongly convex smooth case, we anticipate that $\phi^{2}\geq\phi_{\text{opt}}^{2}$. Without loss of generality, we impose $a_{T}=1$, from which we deduce that $a_{t}=\phi^{2(T-t)}$. We were also able to deduce that $e_{t}=ea_{t+1}$ and that $\alpha_{t}=\alpha a_{t+1}$ for some constants $e,\alpha>0$, yet to be determined.
It remains to study the conditions of Theorem B.4 to find feasible values for $\alpha$ and $e$.
@@
$\begin{cases}a&=(L+\mu)^{2}-4\mu L=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}\phi^{2}=\gamma^{2}(\phi_{opt}^{2}+\varepsilon).\end{cases}$ (7)
It is a simple exercise to verify that $\phi_{opt}=1-\mu\gamma$ if $\gamma\leq\gamma_{opt}$ and $\phi_{opt}=\gamma L-1$ if $\gamma\geq\gamma_{opt}$, which leads us to compute the value of $b$ through a case distinction:
  • if $\gamma\leq\gamma_{opt}$ then $1-\phi_{opt}^{2}=\gamma\mu(2-\gamma\mu)$ and
@@
    $=$ $\gamma\phi_{opt}(L-\mu)+\varepsilon.$
  • if $\gamma\geq\gamma_{opt}$ then $1-\phi_{opt}^{2}=\gamma L(2-\gamma L)$ and
@@
    $=$ $\gamma\phi_{opt}(L-\mu)+\varepsilon.$
In both cases we obtain $b=\gamma\phi_{opt}(L-\mu)+\varepsilon\geq 0$.
@@
$=$ $4\left(\gamma\phi_{opt}(L-\mu)+\varepsilon\right)^{2}-4(L-\mu)^{2}\gamma^{2}(\phi^{2}_{opt}+\varepsilon)$
$=$ $4\gamma^{2}\phi_{opt}^{2}(L-\mu)^{2}+4\varepsilon^{2}+8\varepsilon\gamma\phi_{opt}(L-\mu)-4(L-\mu)^{2}\gamma^{2}\phi^{2}_{opt}-4(L-\mu)^{2}\gamma^{2}\varepsilon$
$=$ $4\left[\varepsilon^{2}+\varepsilon\left(2\gamma\phi_{opt}(L-\mu)-(L-\mu)^{2}\gamma^{2}\right)\right]$
$=$ $4\left[\varepsilon^{2}+\varepsilon\gamma(L-\mu)\left(2\phi_{opt}-(L-\mu)\gamma\right)\right].$
@@
$2\phi_{opt}-(L-\mu)\gamma=\begin{cases}2-\gamma(L+\mu)&\text{ if }\gamma\leq\gamma_{opt}\\ -2+\gamma(L+\mu)&\text{ if }\gamma\geq\gamma_{opt}\end{cases}=|2-\gamma(L+\mu)|=(L+\mu)|\gamma-\gamma_{opt}|,$
such that, with $\delta=|\gamma-\gamma_{opt}|$,
@@
As such, $P$ has real roots. We thus define
@@
$=$ $\frac{\gamma\phi_{opt}(L-\mu)+\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}$
@@
where
$\omega=\frac{\gamma\phi_{opt}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}.$
To conclude the proof, we are now going to verify that this choice of $\alpha$ is feasible, in the sense that it satisfies all the conditions from Theorem B.4. The nonnegativity Condition 1 is satisfied by construction. Condition 2 is satisfied as we set $\rho=0$ and $\alpha_{t}=\beta_{t}$. Condition 3 is equivalent to
@@
We have $\alpha=\omega+\omega_{\varepsilon}\geq\omega$ and $\phi^{2}\geq\phi^{2}_{opt}$, and the above is thus true if
$1\leq\mu L2\omega+\phi^{2}_{opt}$
$\iff$ $1\leq\frac{2\mu L\gamma\phi_{opt}}{L-\mu}+\phi^{2}_{opt}$
$\iff$ $0\leq 2\mu L\gamma\phi_{opt}-(1-\phi^{2}_{opt})(L-\mu).$
@@
  • if $\gamma\leq\gamma_{opt}$, then $\phi_{opt}=1-\gamma\mu$, so the inequality becomes
    00 0 $\leq$ $2\mu L\gamma(1-\gamma\mu)-\gamma\mu(2-\gamma\mu)(L-\mu)$
@@
  • if $\gamma\geq\gamma_{opt}$, then $\phi_{opt}=\gamma L-1$, so the inequality becomes
    00 0 $\leq$ $2\mu L\gamma(\gamma L-1)-\gamma L(2-\gamma L)(L-\mu)$
@@
$2\alpha-\gamma^{2}$ $=$ $2\omega_{\varepsilon}+2\omega-\gamma^{2}=2\omega_{\varepsilon}+\frac{\gamma}{L-\mu}\left(2\phi_{opt}-\gamma(L-\mu)\right)$
@@
Condition 5 is true by the choice of $\alpha$. Finally, Condition 6 holds true provided $02\alpha-\gamma^{2}>$ 0, which is verified through the above computation provided $\varepsilon>0$ or $\delta\neq 0$. In this case,
@@
$=$ $\gamma^{2}\frac{2\gamma\phi_{opt}+2(L-\mu)\omega_{\varepsilon}}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}$
@@
The bound on $e_{T}^{sum}$ is a direct consequence of the fact that $a_{t}$ is a geometric sequence. The upper bound for $e$ comes from the lower bound $\omega_{\varepsilon}\geq\frac{2\varepsilon}{(L-\mu)^{2}}$. ∎
@@
Proposition D.2 (Lyapunov parameters. Strongly convex case, any step-sizes, $L=\mu$).

### v2

@@ A.1.2 Critical Step-Size: $\gamma L=1$
Proposition 7(Bias-optimality. Convex case, large step-sizes)
The bias obtained in the bound of Theorem 4.3 is optimal, in the sense that, for any $x_{0}\in\mathcal{H}$, there exists a convex and smooth problem for which $m=1$ and the inequality is satisfied with equality.
Proof
Consider the function $f(x)=\tfrac{L}{2}\|x\|^{2}$ with minimal value $\inf f=0$ and minimizer $x_{*}=\mathbf{0}$, and take any $x_{0}\in\mathcal{H}$. The iterates $(x_{t})$ of gradient descent are then given by $x_{t}=(1-\gamma L)^{t}x_{0}$, such that
$\frac{1}{T}\sum_{t=0}^{T-1}f(x_{t})-\inf f$ $=\frac{L}{2T}\sum_{t=0}^{T-1}(1-\gamma L)^{2t}\cdot\|x_{0}\|^{2}=\frac{1-(1-\gamma L)^{2T}}{2\gamma(2-\gamma L)T}\cdot\|x_{0}-x_{*}\|^{2}.$
This indeed matches the bias term in Theorem 4.3.
A.1.4 General Step-Size, Suboptimal Bias (Proof of Theorems 4.2 and 4.4)
We now provide a proof for our suboptimal bounds, which cover at the same time the critical and large step-size regimes.
Theorem A.1(Convex case, any step-size, suboptimal bounds)
Let $\gamma L\in(0,2)$ and $\varepsilon\in(0,1)$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
    $\rho=2\gamma(1-\varepsilon)\eta$ where $\eta=1-(\gamma L-1)_{+}$,
  • $\bullet$
    $a_{t}=1$,
  • $\bullet$
    $\alpha_{t}=\gamma/L$,
  • $\bullet$
    $\beta_{t}=\alpha_{t}-\tfrac{\rho}{2L}$,
  • $\bullet$
    $e_{t}=\gamma^{2}\frac{\alpha_{t}+\beta_{t}}{\alpha_{t}+\beta_{t}-\gamma^{2}}$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem 3.1. Moreover, the following simplified bound
$\frac{\bar{e}_{T}}{\rho}\leq\gamma\cdot\frac{\gamma L+\varepsilon\eta}{2(1-\varepsilon)\varepsilon\eta^{2}}$
holds true. In particular, Theorems 4.2 and 4.4 hold true.
Proof
The goal is to find parameters satisfying the sufficient Lyapunov conditions of Theorem 3.1 with a suboptimal value for $\rho$. Therefore we start by defining $\rho=2\gamma(1-\varepsilon)\eta$ for some $\varepsilon\in(0,1)$ and where $\eta=1$ or $2-\gamma L$ depending whether $\gamma L\leq 1$ or $\gamma L\geq 1$. Taking inspiration from our numerical findings, we allow ourselves, in this suboptimal case, to take a sequence $a_{t}$ constantly equal to $1$, which automatically satisfies Condition 3. By Condition 5, we must have $\alpha_{t}=\gamma/L$. We note that Conditions 2 and 4 respectively provide upper and lower bounds on $\beta_{t}$. In view of satisfying Condition 6 with the smallest possible value for $e_{t}$, we aim to pick $\beta_{t}$ to be as large as possible. We therefore saturate the upper bound on $\beta_{t}$ in Condition 2 and take
$\beta_{t}=\alpha_{t}-\frac{\rho}{2L}=\frac{\gamma}{L}-\frac{2\gamma(1-\varepsilon)\eta}{2L}=\frac{\gamma}{L}\left[1-\eta+\eta\varepsilon\right].$
This choice is compatible with Condition 4 as
$\gamma^{2}\leq\alpha_{t}+\beta_{t}\iff\gamma L-1\leq 1-\eta+\eta\varepsilon,$
which is true since
$\gamma L-1\leq(\gamma L-1)_{+}=1-\eta\leq 1-\eta+\eta\varepsilon.$
Finally, the choice of $e_{t}$ is such that Condition 6 is an equality, and hence also satisfied. We note that Condition 1 is verified, as $\beta_{t}\geq 0$ since $\rho\leq 2\gamma$, and the other parameters are naturally nonnegative.
To derive the simplified bound it suffices to compute
$\frac{\bar{e}_{T}}{\rho}=\frac{\gamma^{2}}{\rho}\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta(1-\varepsilon)}=\frac{\gamma}{2(1-\varepsilon)\eta}\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta+\eta\varepsilon}.$
We may simplify the second fraction of this variance term, by considering two cases;
  • $\bullet$
    if $\gamma L\geq 1$, then $\eta=2-\gamma L$ and
    $\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta+\eta\varepsilon}=\frac{2-(2-\gamma L)(1-\varepsilon)}{\eta\varepsilon}=\frac{\gamma L+(2-\gamma L)\varepsilon}{\eta\varepsilon}=\frac{\gamma L+\eta\varepsilon}{\eta\varepsilon}.$
  • $\bullet$
    if $\gamma L\leq 1$, then $\eta=1$ and
    $\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta+\eta\varepsilon}=\frac{1+\varepsilon}{1-\gamma L+\varepsilon}.$
    The above fraction can be seen as a convex function $\phi(\gamma L)$ of $\gamma L\in(0,1)$. As such we can write that $\phi(\gamma L)\leq\phi(0)+\gamma L(\phi(1)-\phi(0))$ where $\phi(0)=1$ and $\phi(1)=\tfrac{1+\varepsilon}{\varepsilon}$, which in our case implies that
    $\frac{1+\varepsilon}{1-\gamma L+\varepsilon}\leq\frac{\gamma L+\varepsilon}{\varepsilon}=\frac{\gamma L+\varepsilon\eta}{\varepsilon\eta}.$
In both cases the second fraction of our variance term is bounded by $\tfrac{\gamma L+\varepsilon\eta}{\varepsilon\eta}$, which gives the desired bound.
Theorems 4.2 and 4.4 now follow by combining Theorem 3.1 and Lemma 7, and noting that $\eta=1$ and $\eta=2-\gamma L$ when $\gamma L=1$ and $\gamma L\geq 1$, respectively.
A.2 Strongly Convex Setting
@@
As in the previous section, we derive parameters that satisfy the sufficient conditions of Theorem 3.1 in order to derive upper bounds for SGD, this time in the strongly convex setting. We again note that all propositions could simply be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
@@
Because it seems impossible to obtain a bound with tight bias term for critical step-sizes (see Sections 7), we provide here a more relaxed bound, where the bias term is $\varepsilon$ away from being tight, and subsequently the variance term remains bounded for every step-size $\gamma L\in(0,2)$.
@@
Proposition 8(Lyapunov parameters. Strongly convex case, any step-size, $L>\mu$)
@@
Let $\gamma L\in(0,2)$, with $L>\mu$. Let $\varepsilon\geq 0$ and $\gamma_{\operatorname{crit}}=\tfrac{2}{\mu+L}$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{\operatorname{crit}}$. Consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
@@
  • $\bullet$
    $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{{\operatorname{opt}}}^{2}+\varepsilon$, with
    $\phi_{{\operatorname{opt}}}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{{\operatorname{opt}}}),$
  • $\bullet$
@@
    $\omega=\frac{\gamma\phi_{{\operatorname{opt}}}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}\quad\text{ with }\quad\delta=|\gamma-\gamma_{\operatorname{crit}}|,$
  • $\bullet$
@@
Then the sufficient Lyapunov conditions of Theorem 3.1 are verified. Moreover, we have
@@
Proof
We start the proof by exploiting some insights from our numerical analysis. First, we observe that we can impose $\rho=0$ and $\alpha_{t}=\beta_{t}$ without changing anything to the empirical best bias term. We moreover infer that ${a_{t}}/{a_{t+1}}\in(0,1)$ is constant, so let us denote this ratio by $\phi^{2}$. Since $\phi_{{\operatorname{opt}}}^{2}$ is the optimal rate for the deterministic gradient descent in the strongly convex smooth case, we anticipate that $\phi^{2}\geq\phi_{\operatorname{opt}}^{2}$. Without loss of generality, we impose $a_{T}=1$, from which we deduce that $a_{t}=\phi^{2(T-t)}$. We were also able to deduce that $e_{t}=ea_{t+1}$ and that $\alpha_{t}=\alpha a_{t+1}$ for some constants $e,\alpha>0$, yet to be determined.
It remains to study the conditions of Theorem 3.1 to find feasible values for $\alpha$ and $e$.
@@
$\begin{cases}a&=(L+\mu)^{2}-4\mu L=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}\phi^{2}=\gamma^{2}(\phi_{{\operatorname{opt}}}^{2}+\varepsilon).\end{cases}$ (7)
It is a simple exercise to verify that $\phi_{{\operatorname{opt}}}=1-\mu\gamma$ if $\gamma\leq\gamma_{\operatorname{crit}}$ and $\phi_{{\operatorname{opt}}}=\gamma L-1$ if $\gamma\geq\gamma_{\operatorname{crit}}$, which leads us to compute the value of $b$ through a case distinction:
  • $\bullet$
    if $\gamma\leq\gamma_{\operatorname{crit}}$ then $1-\phi_{{\operatorname{opt}}}^{2}=\gamma\mu(2-\gamma\mu)$ and
@@
    $=$ $\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon.$
  • $\bullet$
    if $\gamma\geq\gamma_{\operatorname{crit}}$ then $1-\phi_{{\operatorname{opt}}}^{2}=\gamma L(2-\gamma L)$ and
@@
    $=$ $\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon.$
In both cases we obtain $b=\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon\geq 0$.
@@
$=$ $4\left(\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon\right)^{2}-4(L-\mu)^{2}\gamma^{2}(\phi^{2}_{{\operatorname{opt}}}+\varepsilon)$
$=$ $4\gamma^{2}\phi_{{\operatorname{opt}}}^{2}(L-\mu)^{2}+4\varepsilon^{2}+8\varepsilon\gamma\phi_{{\operatorname{opt}}}(L-\mu)-4(L-\mu)^{2}\gamma^{2}\phi^{2}_{{\operatorname{opt}}}-4(L-\mu)^{2}\gamma^{2}\varepsilon$
$=$ $4\left[\varepsilon^{2}+\varepsilon\left(2\gamma\phi_{{\operatorname{opt}}}(L-\mu)-(L-\mu)^{2}\gamma^{2}\right)\right]$
$=$ $4\left[\varepsilon^{2}+\varepsilon\gamma(L-\mu)\left(2\phi_{{\operatorname{opt}}}-(L-\mu)\gamma\right)\right].$
@@
$2\phi_{{\operatorname{opt}}}-(L-\mu)\gamma=\begin{cases}2-\gamma(L+\mu)&\text{ if }\gamma\leq\gamma_{\operatorname{crit}}\\ -2+\gamma(L+\mu)&\text{ if }\gamma\geq\gamma_{\operatorname{crit}}\end{cases}=|2-\gamma(L+\mu)|=(L+\mu)|\gamma-\gamma_{\operatorname{crit}}|.$
If we introduce $\delta\coloneq|\gamma-\gamma_{\operatorname{crit}}|$ then we have the simpler and nonnegative expression
@@
As such, $P$ has real roots. We thus define $\alpha$ to be the largest of the two roots
@@
$=$ $\frac{\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}$
@@
where we noted
$\omega=\frac{\gamma\phi_{{\operatorname{opt}}}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}.$
To conclude the proof, we are now going to verify that this choice of $\alpha$ is feasible, in the sense that it satisfies all the conditions from Theorem 3.1. The nonnegativity Condition 1 is satisfied by construction. Condition 2 is satisfied as we set $\rho=0$ and $\alpha_{t}=\beta_{t}$. Condition 3 is equivalent to
@@
We have $\alpha=\omega+\omega_{\varepsilon}\geq\omega$ and $\phi^{2}\geq\phi^{2}_{{\operatorname{opt}}}$, and the above is thus true if
$1\leq\mu L2\omega+\phi^{2}_{{\operatorname{opt}}}$
$\iff$ $1\leq\frac{2\mu L\gamma\phi_{{\operatorname{opt}}}}{L-\mu}+\phi^{2}_{{\operatorname{opt}}}$
$\iff$ $0\leq 2\mu L\gamma\phi_{{\operatorname{opt}}}-(1-\phi^{2}_{{\operatorname{opt}}})(L-\mu).$
@@
  • $\bullet$
    if $\gamma\leq\gamma_{\operatorname{crit}}$, then $\phi_{{\operatorname{opt}}}=1-\gamma\mu$, so the inequality becomes
    $0$ $\leq$ $2\mu L\gamma(1-\gamma\mu)-\gamma\mu(2-\gamma\mu)(L-\mu)$
@@
  • $\bullet$
    if $\gamma\geq\gamma_{\operatorname{crit}}$, then $\phi_{{\operatorname{opt}}}=\gamma L-1$, so the inequality becomes
    $0$ $\leq$ $2\mu L\gamma(\gamma L-1)-\gamma L(2-\gamma L)(L-\mu)$
@@
$2\alpha-\gamma^{2}$ $=$ $2\omega_{\varepsilon}+2\omega-\gamma^{2}=2\omega_{\varepsilon}+\frac{\gamma}{L-\mu}\left(2\phi_{{\operatorname{opt}}}-\gamma(L-\mu)\right)$
@@
Condition 5 is true by the choice of $\alpha$. Finally, Condition 6 holds true provided $2\alpha-\gamma^{2}>0$, which is verified through the above computation provided $\varepsilon>0$ or $\delta\neq 0$. In this case,
@@
$=$ $\gamma^{2}\frac{2\gamma\phi_{{\operatorname{opt}}}+2(L-\mu)\omega_{\varepsilon}}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}$
@@
The bound on $e_{T}^{sum}$ is a direct consequence of the fact that $a_{t}$ is a geometric sequence. The upper bound for $e$ comes from the lower bound $\omega_{\varepsilon}\geq\frac{2\varepsilon}{(L-\mu)^{2}}$.
@@
Proposition 9(Lyapunov parameters. Strongly convex case, any step-size, $L=\mu$)